# DQ-DETR: DETR with Dynamic Query for Tiny Object Detection

### v3

@@ 1 Introduction
    We design a simple yet accurate categorical counting module to guide the number of object queries effectively. In addition, we enhance the transformer’s visual feature with a density map to improve the positional information of object queries.
@@ 2 Related work
Our proposed DQ-DETR stands out as the first DETR-like model specifically designed to detect tiny objects and dynamically adjust the number of queries to enhance precision in imbalanced datasets.
@@ 3 Method
Figure 1: Overall architecture of our method. (a) Categorical Counting Module, which classifies the number of instances in images into 4 levels. (b) Counting-Guided Feature Enhancement, which refines the encoder’s visual feature with a density map. (c) Dynamic Query Selection, which dynamically adjusts the number of queries and enhances the content and position of queries.
@@ 3.1 Overview
The overall structure of DQ-DETR is shown in Fig. 1. As a DETR-like method, DQ-DETR is an end-to-end detector that contains a CNN backbone, a deformable encoder and decoder [30], and several prediction heads. We further implement a new categorical counting module, a counting-guided feature enhancement module, and dynamic query selection based on DETR’s architecture. Given an input image, we first extract multi-scale features with a CNN backbone and feed them into the transformer encoder to attain visual features. Afterward, our categorical counting module determines how many object queries are used in the transformer decoder, as shown in Fig. 1(a). Besides, we propose a novel counting-guided feature enhancement module, as illustrated in Fig. 1(b), to strengthen the encoder’s visual features with spatial information of tiny objects. Last, the object queries are refined with additional information about the location and size of tiny objects through dynamic query selection, as shown in Fig. 1(c). The following section will describe the proposed Categorical Counting Module, Counting-Guided Feature Enhancement, and Dynamic Query Selection.
@@ 3.3 Categorical Counting Module
We take the largest feature map $S_{1}$ of the EMSV features and generate the density map ${F}_{c}$ through the density extractor. The input feature map $S_{1}$ is sent into a series of dilated convolution layers to acquire a density map ${F}_{c}$, which contains counting-related information. Specifically, dilated convolution layers enlarge the receptive field and capture rich long-range dependency for tiny objects.
@@ 3.4 Counting-Guided Feature Enhancement Module (CGFE)
The EMSV feature is refined using the density map from the categorical counting module through the proposed Counting-Guided Feature Enhancement Module (CGFE) to improve the spatial information of tiny objects. The refined features are then used to enhance the position information of queries.
This module comprises spatial cross-attention and channel attention operations [23].
@@ 3.5 Dynamic Query Selection
Further, we use the intensified multi-scale feature maps ${F}_{t}$ from the previous CGFE module to improve the content ${Q}_{content}$ and position ${Q}_{position}$ of queries. Each layer of ${F}_{t}$ is firstly flattened into pixel level and concatenated together, forming ${F}_{flat}\in\mathbb{R}^{b\times 256\times hw}$. The top-K features will be selected as priors to enhance decoder queries, where $K$ is the number of queries used in the transformer decoder stage. The selection is based on the classification score. We feed ${F}_{flat}$ into an FFN for the object classification task and generate the classification score $\in\mathbb{R}^{b\times m\times hw}$, where m is the number of object classes in the dataset. Consequently, we generate the content and position of queries using the selected top-K features ${F}_{select}$.
@@ 3.6.1 Hungarian Loss
In addition, we use the cross-entropy loss in the categorical counting module to supervise the classification task. Further, the Hungarian loss is also applied as the auxiliary loss for each decoder stage. The overall loss can be denoted as Eq. 8.
@@ 4.2 Implementation Details
4.3 Main Results on AI-TOD-V2
Table 4.2 shows our main results on the AI-TOD-V2 test split. We compare the performances of our DQ-DETR with strong baselines, including both CNN-based and DETR-like methods. All CNN-based methods except YOLOv3 use ResNet50 with feature pyramid network (FPN) [8]. Moreover, since there is no previous research of DETR-like models on tiny object detection, our DQ-DETR is the first DETR-like model that focuses on detecting tiny objects. We re-implement a series of DETR-like models on AI-TOD-V2, and all DETR-like methods except DETR use 5-scale feature maps with deformable attention [30]. For 5-scale feature maps, features are extracted from stages 1, 2, 3, and 4 of the backbone, and add the extra feature by down-sampling the output of stage 4.
@@ 4.3 Main Results on AI-TOD-V2
4.4 Results on VisDrone
Besides AI-TOD-V2, we also conduct experiments on the VisDrone [29] dataset to demonstrate the effectiveness of our model DQ-DETR.
Table 3 shows our results on the VisDrone val split. We compare the performances of our DQ-DETR with other methods. Our proposed DQ-DETR achieves the best result 37.0 AP compared with other state-of-the-art methods, including CNN-based and DETR-like methods. Also, DQ-DETR surpasses the baseline DINO-DETR by 1.2, 2.6, and 1.1 in terms of AP, $\text{AP}_{50}$, $\text{AP}_{75}$.
4.5 Results on COCO
We compare our method, DQ-DETR, with the previous state-of-the-art on the COCO dataset. DQ-DETR yielded slightly lower performance, with an AP of 50.2 compared to 51.3. We believe several factors contributed to these results. Firstly, our experiments were constrained by limited GPU resources, which may have impacted our ability to optimize the training process. Secondly, our method is specifically designed for tiny object detection in scenarios where the number of objects varies significantly across images. This specialized focus may not be fully leveraged in the COCO dataset, which is a general object detection task with a nearly balanced number of objects per image. Therefore, while our method shows potential, it may not perform as expected on general datasets like COCO due to various factors.
@@ 4.6 Ablation Study
Table 6: Evaluation results with different numbers of instances in the images. N indicates the number of instances in the image. We separate the AI-TOD-V2 dataset into 4 classes based on N. All models are trained on the AI-TOD-V2 trainval split and evaluated on test split.

### v5

@@ 1 Introduction
    We propose three components: the categorical counting module, counting-guided feature enhancement, and dynamic query selection. These components significantly enhance performance on tiny objects.
@@ 2 Related work
rmore, while DETR-like methods have improved the formulation of queries, they are constrained in their ability to handle tiny objects and datasets with widely varying numbers of objects. Our proposed DQ-DETR stands out as the first DETR-like model specifically designed to detect tiny objects and dynamically adjust the number of queries to enhance precision in imbalanced datasets.
@@ 3 Method
Figure 1: The overall architecture of our method. (a) Categorical Counting Module, which classifies the number of instances in images into 4 levels. (b) Counting-Guided Feature Enhancement, which refines the encoder’s visual feature with a density map. (c) Dynamic Query Selection, which dynamically adjusts the number of queries and enhances the content and position of queries.
@@ 3.1 Overview
The overall structure of DQ-DETR is shown in Fig. 1. As a DETR-like method, DQ-DETR is an end-to-end detector that contains a CNN backbone, a deformable encoder and decoder [31], and several prediction heads. We further implement a new categorical counting module, a counting-guided feature enhancement module, and dynamic query selection based on DETR’s architecture. Given an input image, we first extract multi-scale features with a CNN backbone and feed them into the transformer encoder to attain visual features. Afterward, our categorical counting module determines how many object queries are used in the transformer decoder, as shown in Fig. 1(a). Besides, we propose a novel counting-guided feature enhancement module, as illustrated in Fig. 1(b), to strengthen the encoder’s visual features with spatial information for tiny objects. Last, the object queries are refined with additional information about the location and size of tiny objects through dynamic query selection, as shown in Fig. 1(c). The following section will describe the proposed Categorical Counting Module, Counting-Guided Feature Enhancement, and Dynamic Query Selection.
@@ 3.3 Categorical Counting Module
We take the largest feature map $S_{1}$ of the EMSV features and generate the density map ${F}_{c}$ through the density extractor. High-resolution features are essential for detecting tiny objects, as they provide a clearer representation of such objects. The input feature map $S_{1}$ is sent into a series of dilated convolution layers to acquire a density map ${F}_{c}$, which contains counting-related information. Specifically, dilated convolution layers enlarge the receptive field and capture rich long-range dependency for tiny objects.
@@ 3.4 Counting-Guided Feature Enhancement Module (CGFE)
The EMSV feature is refined using the density map from the categorical counting module through the proposed Counting-Guided Feature Enhancement Module (CGFE) to improve the spatial information of tiny objects. The refined features are then used to enhance the position information of queries. This module comprises spatial cross-attention and channel attention operations [24].
@@ 3.5 Dynamic Query Selection
Further, we use the intensified multi-scale feature maps ${F}_{t}$ from the previous CGFE module to improve the content ${Q}_{content}$ and position ${Q}_{position}$ of queries. Each layer of ${F}_{t}$ is firstly flattened into pixel level and concatenated together, forming ${F}_{flat}\in\mathbb{R}^{b\times 256\times hw}$. The top-K features are selected as priors to enhance decoder queries, where $K$ is the number of queries used in the transformer decoder stage. The selection is based on the classification score. We feed ${F}_{flat}$ into an FFN for the object classification task and generate the classification score $\in\mathbb{R}^{b\times m\times hw}$, where m is the number of object classes in the dataset. Consequently, we generate the content and position of queries using the selected top-K features ${F}_{select}$.
@@ 3.6.1 Hungarian Loss
In addition, we use the cross-entropy loss in the categorical counting module to supervise the classification task. Further, the Hungarian loss is also applied as the auxiliary loss for each decoder stage. The overall loss can be denoted as:
@@ 4.2 Implementation Details
4.3 Main Results
AI-TOD-V2. Table 4.2 shows our main results on the AI-TOD-V2 test split. We compare the performances of our DQ-DETR with strong baselines, including both CNN-based and DETR-like methods. All CNN-based methods except YOLOv3 use ResNet50 with feature pyramid network (FPN) [8]. Moreover, since there is no previous research on DETR-like models for tiny object detection, our DQ-DETR is the first DETR-like model that focuses on detecting tiny objects. We re-implement a series of DETR-like models on AI-TOD-V2, and all DETR-like methods except DETR use 5-scale feature maps with deformable attention [31]. For 5-scale feature maps, features are extracted from stages 1, 2, 3, and 4 of the backbone, and add the extra feature by down-sampling the output of stage 4.
@@ 4.3 Main Results
VisDrone. We also conduct experiments on the VisDrone [30] dataset to demonstrate the effectiveness of our model DQ-DETR. Table 3 shows our results on the VisDrone val split. We compare the performances of our DQ-DETR with other methods. Our proposed DQ-DETR achieves the best result 37.0 AP compared with other state-of-the-art methods, including CNN-based and DETR-like methods. Also, DQ-DETR surpasses the baseline DINO-DETR by 1.2, 2.6, and 1.1 in terms of AP, $\text{AP}_{50}$, $\text{AP}_{75}$.
COCO.       We compare our method, DQ-DETR, with the previous state-of-the-art on the COCO dataset. DQ-DETR yielded slightly lower performance, with an AP of 50.2 compared to 51.3. We believe several factors contributed to these results. Firstly, our experiments were constrained by limited GPU resources, which may have impacted our ability to optimize the training process. Secondly, our method is specifically designed for tiny object detection in scenarios where the number of objects varies significantly across images. This specialized focus may not be fully leveraged in the COCO dataset, which is a general object detection task with a nearly balanced number of objects per image. Therefore, while our method shows potential, it may not perform as expected on general datasets like COCO due to various factors.
@@ 4.4 Ablation Study
Table 6: Evaluation results for varying instance counts. N indicates the number of instances in the image. We separate the AI-TOD-V2 dataset into 4 classes based on N. All models are trained on the AI-TOD-V2 trainval split and evaluated on test split.
@@ Acknowledgment
Appendix
Appendix 0.A Quantitative Results
0.A.1 FP/FN Under Different Density Situation
In Table A1, we explore our DQ-DETR’s performance under different density situations, including sparse and dense images. We classify images with less than 100 objects as sparse and images with over 900 objects as dense. LRP FP and LRP FN [14] are used as the evaluation metric. Unlike AP metrics, a lower LRP value implies better performance. Previous DETR-like models apply a fixed number of object queries, while our DQ-DETR uses a dynamic number of queries depending on the object’s density in the picture.
DINO-DETR uses a fixed number of 900 queries for detection, no matter whether in dense or sparse situations. The number of queries exceeds the number of objects in the sparse image and hence introduces many underlying false positive samples (FP). In contrast, for dense images, the number of queries DINO-DETR uses is far less than the number of objects in images, which is beyond the detection capability of DINO-DETR, leaving lots of instances undetected (FN) and causes a large LRP FN score. Our proposed DQ-DETR dynamically adjusts the number of queries used for detection, resulting in fewer FP in sparse images and fewer FN in dense images.
Table A1: LRP FP and LRP FN score under different density situations in AI-TOD-V2. DINO-DETR is compared as the baseline.
{tblr}
width = column3 = c, column4 = c, column2 = c, cell21 = r=2, cell41 = r=2, vline2 = 1-5, vline3 = 1-5, hline1-2 = -, hline4,6 = -, Method Situation LRP FP LRP FN
DINO-DETR Sparse 29.4 40.7
Dense 36.8 75.1
DQ-DETR Sparse 25.7 36.4
Dense 35.4 51.5
0.A.2 Ablation of Categorical Counting Module
In the categorical counting module, we categorize the number of objects $N$ per image into four levels, which are $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $N>500$. We selected the numbers 10, 100, and 500 based on the AI-TOD-V2 dataset’s characteristics, i.e., the mean and standard deviation of the number of instances $N$ per image. Further, we only classify the number of objects $N$ into four levels due to the long-tail distribution of the training samples. For the $N>500$ situation, there are only 46 training images in this situation, which is much fewer than other cases and leads to a poor classification performance with only 56.6% accuracy. Although there are at most 2267 instances per image in the AI-TOD-V2 dataset, the long-tail distribution of the training samples restricts us from classifying the number of instances $N$ per image in a more detailed manner.
Table A2 and Table A3 demonstrate the detection performance and the accuracy of the classification task in our categorical counting module. We can observe that if we classify the number of objects $N$ into more classes, e.g., 5 classes, AP drops 1.4 compared to the 4-class scenario. That is because the poor classification results from the categorical counting module will directly affect the number of object queries used for detection and the inappropriate number of queries might harm the detection performance. In the 5-class classification scenario, while the total classification accuracy maintains 93.8%, the accuracy in the $500<$ N $\leq 900$, and $N>900$ situations are only 37.1% and 57.4%. Since there are only a few training images, the categorical counting module doesn’t perform well in these two situations and further impacts the detection performance. Hence, we only categorize the number of objects $N$ per image into four levels without partitioning the $N>500$ situation into more detailed settings.
Table A2: Ablation of categorical counting module. DINO-DETR is compared as the baseline.
{tblr}
columneven = c, column3 = c, column5 = c, vline2 = -, hline1-2,5 = -, Method AP $\text{AP}_{vt}$ $\text{AP}_{t}$ $\text{AP}_{s}$ $\text{AP}_{m}$
Baseline 25.9 12.7 25.3 32.0 39.7
Classification (4cls) 30.2 15.3 30.5 36.5 44.6
Classification (5cls) 28.8 14.3 29.2 34.1 43.1
Table A3: The classification accuracy of our categorical counting module with different numbers of classes.
{tblr}
width = columneven = c, column3 = c, vline2-4 = -, hline1-2,7-8 = -, #Objects in image Accuracy(%) @ 4cls Accuracy(%) @ 5cls #Sample
N $\leq 10$ 97.7 97.5 8674
$10<$ N $\leq 100$ 90.5 89.3 4393
$100<$ N $\leq 500$ 86.5 83.2 905
$500<$ N $\leq 900$ 56.5 37.1 35
$900<$ N - 54.4 11
Total 94.6 93.8 14018
0.A.3 Categorical counting module (CCM) for different datasets.
Since the characteristics of different datasets may vary a lot, it is vital to tailor our categorical counting module to the dataset property and determine how many queries should be used for object detection. In the categorical counting module for AI-TOD-V2, we estimate the counting number $N$, i.e., the number of instances per image, by a classification head and categorize them into four levels, which are $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $N>500$. It is worth noting that the hyperparameters 10, 100, and 500 are tailored for AI-TOD-V2. For other datasets, we recommend adjusting the hyperparameters used in the CCM through a logical process that considers the mean and variance of the objects per image in the dataset. This approach can reduce the need to manually design the CCM for different datasets.
Algorithm 1 Pseudo Code for Categorical Counting Module
1:function Categorical-Counting($features$)
2:     $mean\leftarrow\text{Mean}(dataset)$
3:     $var\leftarrow\text{Variance}(dataset)$
4:     $GT\leftarrow\text{Predict}(features)$
5:     $class1\leftarrow\{GT<mean-var\}$
6:     $class2\leftarrow\{mean-var\leq GT<mean\}$
7:     $class3\leftarrow\{mean\leq GT<mean+var\}$
8:     $class4\leftarrow\{GT\geq mean+var\}$
9:end function
Figure B1: Visualization of detection results and feature maps. The green, red, and blue boxes represent TP, FP, and FN, respectively.
Appendix 0.B Visualization
Fig. B1 presents the detection results of our DQ-DETR alongside Deformable-DETR. By selecting an appropriate number of object queries, DQ-DETR effectively detects most of the tiny objects in dense scenes, demonstrating superior performance in capturing fine details. Conversely, Deformable-DETR suffers from an insufficient number of object queries, leading to a higher rate of undetected objects (false negatives). In sparse scenes, Deformable-DETR further struggles due to the use of excessive object queries with unrefined positional information, resulting in an increased number of false positives in the detection outcomes.